Operational issues:
MGHPCC site level cooling issue for ~1 hour caused NESE Ceph to be offline for about 1 hour.
GPFS disk errors in the system pool needed evacuation. Reduced SGE while this was in process.
88 worker nodes installed and tested, all but 1 rack in production.
Rack of contributed nodes arrived.
Hard to get 100Gb optics from DELL.
New NESE Ceph storage equipment arrived except for Cisco switches, expected in August following long delays.
NESE Tape commissioning: 50PB pool being reformatted today re: IBM firmware issue.
Working with NESE team on NESE Tape expansion.
Quarterly reports in. Hardware spreadsheet updated.
Run 3 software prep: Working on perfsonar & ipv6. OSG 3.6 to follow.